Johnson Farms is planning to install one of two mechanical devices to reduce costs. Both cost $1000 and last five years with no salvage value. Device A can be expected to result in $300 savings annually. Device B will provide cost saving of $400 the first year but will decline by $50 annually, making the second year saving to $350 and third year savings $300 and so forth, with interest at 7% per year, calculate PW of each device. Which device should the firm purchase? Why?
Answer #1

MARR, i

8%

p.a.

device fixed cost

$1,000


useful life, n

                5

years

Both the devices here cost $1000, and the decision criterion will be to choose the device that maximizes the present worth of the benefits.

The Present worth of cost of each device is $1000.




Present worth of benefits:

Device A = 400(P/A,8%,5)

(here $400 is the saving from the device every year)

Using standard COMPOUND INTEREST TABLES: (i=8%,n=5)

PW of benefits for A = 400(3.993)


$1,597.20


PW of Device A = PW of benefits for A - PW of cost


$597.20





Present worth of benefits:

Device B = 400(P/A,8%,5)-50(P/G,8%,5)

Using standard COMPOUND INTEREST TABLES: (i=8%,n=5)

PW of benefits for B = 400 (3.993) - 50(7.372)

PW of benefits for B =

$1,228.60


PW of Device B = PW of benefits for B - PW of cost


$228.60





PW of each device:



PW


Device A

$597.20


Device B

$228.60





Hence, device A should be chosen as it has higher present worth.


answered by: ANURANJAN SARSAM